OpenStudy (anonymous):

After isolating the absolute value in the equation -2|5n + 10| = -30, you would see that there is “no solution” for this problem.

OpenStudy (anonymous):

i don't think so

OpenStudy (anonymous):

first divide both sides by \(-2\) to get \(|5n+10|=15\) then you should get two solutions

OpenStudy (anonymous):

one when you solve \(5n+10=15\) and the other when you solve \(5n+10=-15\)

OpenStudy (anonymous):

the answer is 8